how to solve the question
the point 3x+y+4=0 which is equidstant from(-5,6)and(3,2)

Grade:

1 Answers

ruchi yadav
10 years ago
let point is (p,q)

so from equation of line 3p+q+4 = 0 (equation 1 )

and now euidistant from (-5,6) and (3,2)

(p+5)2 + (q-6)2 = (p-3)2 +(q-2)2

two equation and two unknown.
p = -5/3
q = 1



Thanks,
Ruchi
Askiitians faculty